Hi!
Find a polynomial with integer coefficients and a leading coefficient of one that satisfies the given conditions.
Q has degree 3, and zeros -9 and 1 + i
thanks for your help!

Answer by edjones(8007)

x=-9, x=1+i, x=1-i
(x+9)(x-(1+i)(x-(1-i)
x^3+7x^2-16x+18
.
Ed